The sum of the third and twice the first of three consecutive odd integers is 49. Find all three consecutive odd integers and the equation.: The sum of the third and twice the first of three consecutive odd integers is 49. Find all three consecutive odd integers and the equation.
Answer by stanbon(18754)

The sum of the third and twice the first of three consecutive odd integers is 49. Find all three consecutive odd integers and the equation.
----------
1st: 2x-1
2nd: 2x+1
3rd: 2x+3
---------
EQUATION:
2(2x-1) + 2x+3 =49
4x-2 + 2x+3 = 49
6x+1 = 49
6x = 48
x = 8
------------------
1st: 2x-1 = 15
2nd: 17
3rd: 19
===============
Cheers,
Stan H.